what 2 digit number is equal to twice the product of it's digits?
Answer by Fombitz(32388)

Let the number be
A*10+B where A and B are between 0 and 9
A*10+B=2(AB)
10+(B/A)=2B
B and A must be multiples of each other for this equation to work.
(2,4) and (3,6) are the only digits that are multiples.
10+(4/2)=2(4)
12=8
No, that doesn't work.
10+(6/3)=2(6)
12=12
Yes, that works.
The number is 36.
36=2*(3*6)=36
.
.
.
The other possible solution is 00, because
00=2(0*0)
although I'm not sure that 0 is really a 2 digit number.